OpenStudy (anonymous):

(50 - 75) ÷ 52 - 7(4)find the value

OpenStudy (mimi_x3):

BEDMAS B=Brackets E=Exponents D/M=division/multiplication A/S= Addition/Subtraction So lets first do the brackets (50-75)=-25 -25/52-7(4) Now lets multiply 7(4)=28 So we have -25/52-28 Its gonna be an ugly number -.- Divide -25 by 52 and then subtract 28